# A Step Toward World Models: A Survey on Robotic Manipulation

### v1

@@ II-D Revisiting Modern AI Models Through the Lens of World Modeling
The rapid progress of large-scale artificial intelligence models has blurred the boundaries between traditional task-specific learning and general world modeling. Although many of these models are not explicitly designed as world models, they exhibit key characteristics of world modeling, such as learning structured representations of reality, reasoning about causality, and predicting or generating plausible future states. Revisiting these modern models through the lens of world modeling provides valuable insights into how intelligence emerges from data, embodiment, and multimodal integration. This perspective helps clarify which components of contemporary architectures, such as large language models (LLMs), vision-language models (VLMs), vision-language-action models (VLAs) and video generation models, implicitly capture aspects of the world and how they contribute to the broader goal of general-purpose world understanding.
II-D1 LLMs, VLMs & VLA
@@
The strong reasoning capabilities and next-token prediction mechanism of Large Language Models (LLMs) make them natural foundations for constructing world models, as they capture sequential dependencies, causal relationships, and abstract dynamics. When equipped with auxiliary modules such as value functions [62] or modality-specific encoders [72, 54, 60], LLMs can achieve a more comprehensive understanding of the environment. Moving beyond the language-centric paradigm, Vision–Language Models (VLMs) focus on the joint modeling of multiple modalities, providing a perceptually grounded understanding of the world [83, 84]. Furthermore, an increasing number of studies have explored augmenting VLMs with low-level action-generation capabilities, thereby transforming them into Vision–Language–Action (VLA) models [16, 85, 44] that bridge perception, reasoning, and control. In addition to action generation, there are also many work that enable additional visual prediction [78, 79].
@@
From the above discussion, the designs and functions of LLMs, VLMs, and VLAs align with the spirit of world models, as they aim to represent and reason about world dynamics. Therefore, these models should not be excluded from the broader conceptual scope of world modeling. However, solely relying on LLMs, VLMs, or VLAs often constrains a system’s capacity for long-horizon prediction, reasoning, and imagination, all of which are essential for modeling dynamic and interactive environments. Recent studies have thus begun to integrate these models into architectures that explicitly function as world models, such as the JEPA framework [27], Dreamer-style frameworks [86], positioning them as core mechanisms for capturing temporal and causal dynamics.
II-D2 Video Generation Models.
@@
Video generation models primarily aim to produce visually realistic and temporally coherent sequences, which implicitly rely on learning the underlying dynamics of the environment. They can operate on diverse modalities, including language, visual data, and action inputs, allowing them to access environmental context and imagine future scenes. These characteristics position video generation models as a form of world modeling. Indeed, many recent world models adopt video generation as their core mechanism [5, 6, 7], enabling the prediction of future states encompassing observations, actions, and environmental changes. However, most video generation models focus on observation-level prediction and may lack interpretable internal representations of the world.
@@
III Overview of the World Model
@@ III-A Paradigms
Building on the previous review of current models, contemporary architectures for capturing world dynamics can be broadly stratified along a methodological spectrum: implicit world modeling (e.g., LLMs, VLMs, and VLAs) [27, 87, 41, 85], latent dynamics modeling [14, 12, 13, 15], and video generation paradigms [6, 5, 7, 67], each targeting distinct representational granularities and predictive mechanisms.
Figure 2: A visualization of LLM-based world models [54].
@@ III-A1 Implicit World Modeling
Representative models include LLMs, VLMs, and VLAs, which offer distinct advantages in semantic grounding, generalization, and interpretability [62, 72, 54, 60, 84]. An illustration of these models is shown in Fig. 2. At the same time, these models can be integrated into broader world-modeling architectures to capture temporal dependencies and enable long-horizon prediction [16, 85, 44]. Detailed discussions of these models are provided in Sections II-D1 and IV-A1.
@@ III-A2 Latent Dynamics Modeling
Latent dynamics models typically encode high-dimensional observations into compact latent states through a variational autoencoder (VAE) or encoder network, and employ recurrent or transformation modules (e.g., RNNs or Transformers) to predict the temporal evolution of these latent representations [12, 13, 14, 15]. This architecture is characterized by latent-space imagination and task-oriented optimization over visual granularity, facilitating long-horizon learning by forecasting future states without the need for pixel-level reconstruction.
(a) Dynamics learning from experience
(b) Behavior learning in imagination
Figure 3: A visualization of Dreamer architecture [12], which encodes visual representations into latent states through recurrent estimation.
Recurrent State-Space Model (RSSM) [29] resembles the structure of a partially observable Markov decision process. Its learning framework consists of three main components: an encoder, a decoder, and a dynamics model. The encoder network fuses sensory inputs (observations) $o$ together into the stochastic representations $z$. The dynamics model learns to predict the sequence of stochastic representations by using its recurrent state $s$. The decoder reconstructs sensory inputs to provide a rich signal for learning representations and enables human inspection of model predictions, but is not needed while learning behaviors from latent rollouts. Specifically, at time step $t$, let the image observation be $o_{t}$, the action vectors $a_{t}$ and the reward $r_{t}$. RSSM can be formulated as the generative process of the images and rewards conditioned a hidden state sequence $s_{t}$:
@@
PlaNet [29] first demonstrates the effectiveness of learning dynamics in a latent space. The Dreamer family of models (a visualization is shown in Fig. 3) [12, 13, 14, 15] further verify this paradigm and establish a representative framework that reduces reliance on real-world data by performing imagination directly in latent space. Dreamer enables policy learning through imagined trajectories without continuous interaction with the environment, allowing agents to simulate multi-step consequences of actions and generalize to new states, objects, and environments.
Figure 4: A visualization of Joint Embedding-Action-Prediction (JEPA) Architecture [11], where self-supervised learning is used to learn the future world state representations.
@@
While sharing the objective of learning predictive world-state representations, Joint-Embedding Predictive Architecture (JEPA) [11, 27] and RSSM diverge fundamentally in their learning mechanisms. RSSM relies on generative reconstruction of observations to model latent dynamics, whereas JEPA (a visualization is shown in Fig. 4) employs self-supervised predictive coding in embedding spaces—directly forecasting future state representations without decoding to raw sensory inputs. This paradigm eliminates the computational cost of pixel-level reconstruction but necessitates powerful hierarchical encoders to compress sufficient environmental information into abstract embeddings, creating an implicit information bottleneck that demands careful architectural balancing to preserve task-relevant features. Under the JEPA framework, Assran et al. [77] combine pre-trained video models with an action-conditioned predictor to autoregressively predict future states and actions.
@@
The MuZero series [88, 89, 90] represent another form of latent-dynamics-based world modeling. Instead of modeling the complete environment dynamics, MuZero predicts only future quantities directly relevant to planning, such as rewards, values, and policies, given the complexity of real-world environments, and employs a tree-based search algorithm [91] to select optimal actions.
III-A3 Video Generation.
Figure 5: An illustration of video-geneation based world models [57]. World model serve as the core component, modelling the world dynamics and enabling action planning and generation.
@@
Video-based generative models are powerful tools for capturing environmental dynamics and predicting future scenes. These models operate directly on high-dimensional raw observations, such as RGB images, depth maps, or force fields [6, 23, 19, 20, 92, 72, 93, 66, 94], treating the environment as a sequence of frames. By generating future scenes, they can support a wide range of applications, including visual planning, simulation, and action generation [63, 64, 6, 7, 67]. Moreover, they can leverage large-scale pre-training to enhance generalization and improve sample efficiency [67, 23, 67, 9, 38]. Depending on the input modality, world models can be constructed using action-conditioned video prediction models [6], text-to-video models [5, 19, 38, 66], or trajectory-to-video models [43, 58].
@@
There are several architectural families of video-based world models. Diffusion-based world models generate videos by progressively denoising random noise through multiple iterative steps. Representative examples include U-Net-based models [95, 70] and diffusion transformer (DiT)-based architectures [25, 96, 43, 97, 10]. Autoregressive world models, in contrast, predict the next token or frame conditioned on previously generated ones, effectively modeling temporal dependencies in the sequence [57, 6, 72, 8, 26, 20, 58]. Other architectures include variational autoencoder (VAE)-based models [20] and convolutional LSTMs [63, 64].
@@
Autoregressive-based world models generate each step conditioned on previous outputs, allowing them to predict sequences of arbitrary length and making them well-suited for long-horizon predictions. However, they often suffer from error accumulation over extended sequences [10] and may struggle to represent highly multi-modal distributions. In contrast, diffusion-based models generate samples through an iterative denoising process, enabling them to model complex, multi-modal distributions and produce globally coherent sequences. This iterative refinement also makes diffusion models more robust to individual prediction errors, resulting in better performance on tasks requiring long-horizon consistency or high-quality generative outputs. On the downside, diffusion models are computationally intensive and slower during inference, and adapting them to sequential prediction requires careful conditioning. Overall, autoregressive world models tend to excel in scenarios demanding speed and accurate short-term predictions, whereas diffusion models are more suitable for tasks involving long-horizon, multi-modal, or high-dimensional outputs where maintaining global coherence is crucial.
@@
Compared with implicit world models and latent-space world models, video generation models provide more detailed visual predictions but at a higher computational cost, lower generation speed and sample efficiency. In addition, action predictions are only proved to be align with visual future generation [73], as visual data contain relevant information to actions.
@@ III-B1 Flat architecture
Most current methods adopt flat architectures [21, 25, 26, 20, 92, 72, 93, 66], which face critical limitations. They lack structured representations of the environment, resulting in poor handling of multi-scale dynamics, limited long-horizon prediction, error accumulation, and reduced generalization. Specifically, in robotic manipulation, placing fragile objects requires the robot to react instantly to unexpected slips while simultaneously planning the sequence of pick-and-place actions to achieve the overall goal. Many tasks further involve long-term objectives that must be completed through sequential subgoals and temporally extended actions. For example, assembling a piece of furniture requires picking up components, aligning and attaching them correctly, and tightening screws for each part. Moreover, operating at a single level of abstraction causes small prediction errors to compound over time, degrading performance in long-horizon tasks. Finally, flat architectures fail to extract high-level abstractions, limiting transferability across tasks and environments.
III-B2 Hierarchical architecture.
Several studies have begun to explore and design hierarchical world models, in which lower-level modules handle intermediate reactions and short-term predictions, while higher-level components are responsible for long-term planning and abstraction. Lecun et al. [11] hypothesize a hierarchical JEPA architecture, where low-level and high-level representations are learned for short- and long-term predictions, respectively. Gumbsch et al. [56] propose an RSSM-based hierarchical world model, where the low-level module captures immediate dynamics for reactive control, and the high-level module models abstract temporal patterns for strategic planning. Björck et al. [44] introduce a dual-system architecture in which System 2 interprets the environment and task goals, while System 1 generates continuous motor commands in real time. Similarly, Wang et al. [45] design a dual-level world model consisting of an RSSM-based System 1 (RSSM-S1) and a logic-integrated neural network System 2 (LINN-S2). The inter-system feedback mechanism ensures that predicted sequences comply with domain-specific logical rules: LINN-S2 constrains RSSM-S1’s predictions, while RSSM-S1 updates LINN-S2 based on new observations, enabling dynamic adaptation. Wang et al. [98] further employ System 2 for value-guided high-level planning by estimating state-action values and selecting optimal actions, while System 1 executes real-time motions via cascaded action denoising.
@@ III-C1 Dimensionality of the World
In designing world models, the dimensionality of the environment plays a critical role, shaping how effectively a model captures spatial structures, temporal evolution, and causal dynamics.
Some works operate purely in 2D pixel space [19, 20, 92, 72, 93, 66], capturing visual appearance and short-term dynamics but ignoring the real-world geometry. While 2D pixel-space models [19] capturing visual appearance and short-term dynamics but lacking geometric awareness of real-world structure. This limitation motivates the development of 3D-aware architectures. To incorporate geometric understanding of the 3D world, Bu et al. [22, 70, 7] construct world models based on RGB-D data, while others extract richer 3D cues such as scene flow [21], motion fields [69] and 3D point clouds with associated language descriptions [41], enabling more comprehensive modeling of 3D world dynamics. Additionally, Lu et al. [34] leverage 3D Gaussian Splatting, Diffusion Transformers, and 3D Gaussian Variational Autoencoder to extract 3D representations from RGB observations. Zhang et al. [60] incorporate depth estimation to enhance the understanding of 3D worlds. In addition to geometric structure, temporal dynamics are incorporated to construct 4D world models that jointly capture spatial and temporal evolution. For example, Zhu et al. [23] synthesize 4D data from RGB-D videos by estimating depth and camera pose. Zhen et al. [24] leverage the pre-trained 3D VAE [99] to encode RGB, depth, and normal videos and sum them together, while Huang et al. [8] employ 4D Gaussian splatting to model spatiotemporal dynamics in robotic environments.

### v2

@@ II-D Revisiting Modern AI Models Through the Lens of World Modeling
The rapid progress of large-scale artificial intelligence models has blurred the boundaries between traditional task-specific learning and general world modeling. Although many of these models are not explicitly designed as world models, they exhibit key characteristics of world modelimodelling as learning structured representations of reality, reasoning about causality, and predicting or generating plausible future states. Revisiting these modern models through the lens of world modeling provides valuable insights into how intelligence emerges from data, embodiment, and multimodal integration. This perspective helps clarify which components of contemporary architectures, such as large language models (LLMs), vision-language models (VLMs), vision-language-action models (VLAs) and video generation models, implicitly capture aspects of the world and how they contribute to the broader goal of general-purpose world understanding.
II-D1 LLM, VLM & VLA
@@
The strong reasoning capabilities and next-token prediction mechanism of LLMs make them natural foundations for constructing world models, as they capture sequential dependencies, causal relationships, and abstract dynamics. When equipped with auxiliary modules such as value functions [ahn2022can] or modality-specific encoders [xiang2024pandora, driess2023palm, zhang2025dreamvla], LLMs can achieve a more comprehensive understanding of the environment. Moving beyond the language-centric paradigm, VLMs focus on the joint modeling of multiple modalities, providing a perceptually grounded understanding of the world [hu2023look, zhao2024vlmpc]. Furthermore, an increasing number of studies have explored augmenting VLMs with low-level action-generation capabilities, thereby transforming them into VLA [zitkovich2023rt, hong20233d, bjorck2025gr00t] that bridge perception, reasoning, and control. In addition to action generation, there are also many work that enable additional visual prediction [zhang2025up, zhao2025cot].
@@
From the above discussion, the designs and functions of LLMs, VLMs, and VLAs align with the spirit of world models, as they aim to represent and reason about world dynamics. Therefore, these models should not be excluded from the broader conceptual scope of world modeling. However, solely relying on LLMs, VLMs, or VLAs often constrains a system’s capacity for long-horizon prediction, reasoning, and imagination, all of which are essential for modeling dynamic and interactive environments. Recent studies have thus begun to integrate these models into architectures that explicitly function as world models, such as the JEPA framework [chen2025egoagent], Dreamer-style frameworks [wang2025founder], positioning them as core mechanisms for capturing temporal and causal dynamics.
II-D2 Video Generation Models
@@
Video generation models primarily aim to produce visually realistic and temporally coherent sequences, which implicitly rely on learning the underlying dynamics of the environment. They can operate on diverse modalities, including language, visual data, and action inputs, allowing them to access environmental context and imagine future scenes. These characteristics position video generation models as a form of world modeling. Indeed, many recent world models adopt video generation as their core mechanism [du2023video, wu2024ivideogpt, zhang2025combo], enabling the prediction of future states encompassing observations, actions, and environmental changes. However, most video generation models focus on observation-level prediction and may suffer from issues such as a lack of interpretable internal representations and limited causal understanding of the world.
@@
III Overview of World Models
@@ III-A Paradigms
Building on the previous review of current models, contemporary architectures for capturing world dynamics can be broadly stratified along a methodological spectrum: implicit world modeling (e.g., LLMs, VLMs, and VLAs) [chen2025egoagent, team2023internlm, zhen20243d, hong20233d], latent dynamics modeling [wu2023daydreamer, hafnerdream, hafner2021mastering, hafner2023mastering], and video generation paradigms [wu2024ivideogpt, du2023video, zhang2025combo, rigter2025avid], each targeting distinct representational granularities and predictive mechanisms.
Figure 3: An overview of world models. Implicit world models map observations and instructions directly to actions without explicitly modeling environmental dynamics. Latent-dynamics world models capture the evolution of environment dynamics within a latent space, while video-generation-based world models predict future visual scenes.
@@ III-A1 Implicit World Modeling
Implicit World Modeling refers to the paradigm in which an agent learns to map sensory inputs (e.g., vision and language) directly to actions without constructing an explicit, disentangled representation of the environment’s states or dynamics. Representative models include LLMs, VLMs, and VLAs, which offer distinct advantages in semantic grounding, generalization, and interpretability [ahn2022can, xiang2024pandora, driess2023palm, zhang2025dreamvla, zhao2024vlmpc]. An illustration of these models is shown in Fig. 3. At the same time, these models can be integrated into broader world-modeling architectures to capture temporal dependencies and enable long-horizon prediction [zitkovich2023rt, hong20233d, bjorck2025gr00t]. Detailed discussions are provided in Sections II-D1 and IV-A1.
@@ III-A2 Latent Dynamics Modeling
Latent dynamics models typically encode high-dimensional observations into compact latent states through a variational autoencoder or encoder network, and employ recurrent or transformation modules (e.g., RNNs or Transformers) to predict the temporal evolution of these latent representations [hafnerdream, hafner2021mastering, wu2023daydreamer, hafner2023mastering]. This architecture is characterized by latent-space imagination and task-oriented optimization over visual granularity, facilitating long-horizon learning by forecasting future states without the need for pixel-level reconstruction.
Recurrent State-Space Model (RSSM) [hafner2019learning] resembles the structure of a partially observable Markov decision process. Its learning framework consists of three main components: an encoder, a decoder, and a dynamics model. The encoder network fuses sensory inputs (observations) $o$ together into the stochastic representations $z$. The dynamics model learns to predict the sequence of stochastic representations by using its recurrent state $s$. The decoder reconstructs sensory inputs to provide a rich signal for learning representations and enables human inspection of model predictions, but is not needed while learning behaviors from latent rollouts. Specifically, at time step $t$, let the image observation be $o_{t}$, the action vectors $a_{t}$ and the reward $r_{t}$. RSSM can be formulated as the generative process of the images and rewards conditioned a hidden state sequence $s_{t}$:
@@
PlaNet [hafner2019learning] first demonstrates the effectiveness of learning dynamics in a latent space. The Dreamer family of models (a visualization is shown in Fig. 3) [hafnerdream, hafner2021mastering, wu2023daydreamer, hafner2023mastering] further verify this paradigm and establish a representative framework that reduces reliance on real-world data by performing imagination directly in latent space. Dreamer enables policy learning through imagined trajectories without continuous interaction with the environment, allowing agents to simulate multi-step consequences of actions and generalize to new states, objects, and environments.
@@
While sharing the objective of learning predictive world-state representations, Joint-Embedding Predictive Architecture (JEPA) [lecun2022path, chen2025egoagent] and RSSM diverge fundamentally in their learning mechanisms. RSSM relies on generative reconstruction of observations to model latent dynamics, whereas JEPA (a visualization is shown in Fig. 3) employs self-supervised predictive coding in embedding spaces, directly forecasting future state representations without decoding to raw sensory inputs. This paradigm eliminates the computational cost of pixel-level reconstruction but necessitates powerful hierarchical encoders to compress sufficient environmental information into abstract embeddings. As a result, it introduces an implicit information bottleneck that demands careful architectural balancing to preserve task-relevant features. Under the JEPA framework, Assran et al. [assran2025v] combine pre-trained video models with an action-conditioned predictor to autoregressively predict future states and actions.
@@
The MuZero series [schrittwieser2020mastering, ye2021mastering, wang2024efficientzero] represent another form of latent-dynamics-based world modeling. Instead of modeling the complete environment dynamics, MuZero predicts only future quantities directly relevant to planning, such as rewards, values, and policies, given the complexity of real-world environments, and employs a tree-based search algorithm [silver2018general] to select optimal actions.
III-A3 Video Generation
@@
Video-based generative models are powerful tools for capturing environmental dynamics and predicting future scenes. These models operate directly on high-dimensional raw observations, such as RGB images, depth maps, or force fields, treating the environment as a sequence of frames [wu2024ivideogpt, team2025aether, yang2023learning, bruce2024genie, brooks2024video, xiang2024pandora, zheng2024open, zhou2024robodreamer, ali2025humanoid]. By generating future scenes, they can support a wide range of applications, including simulation, action prediction, and visual planning [finn2017deep, ebert2018visual, wu2024ivideogpt, zhang2025combo, rigter2025avid]. Moreover, they can leverage large-scale pre-training to enhance generalization and improve sample efficiency [rigter2025avid, team2025aether, rigter2025avid, wang2025language, jang2025dreamgen]. Depending on the input modality, world models can be constructed using action-conditioned video prediction models [wu2024ivideogpt], text-to-video models [du2023video, yang2023learning, jang2025dreamgen, zhou2024robodreamer], or trajectory-to-video models [zhu2025irasim, cheang2024gr].
@@
There are several architectural families of video-based world models. Diffusion-based world models generate videos by progressively denoising random noise through multiple iterative steps. Representative examples include U-Net-based models [ho2022video, ko2024learning] and diffusion transformer (DiT)-based architectures [ferraro2025focus, agarwal2025cosmos, zhu2025irasim, wan2025wan, yang2025roboenvision]. Autoregressive world models, in contrast, predict the next token or frame conditioned on previously generated ones, effectively modeling temporal dependencies in the sequence [liao2025genie, wu2024ivideogpt, xiang2024pandora, huang2025enerverse, villar2025playslot, bruce2024genie, cheang2024gr]. Other architectures include variational autoencoder-based models [bruce2024genie] and convolutional LSTMs [finn2017deep, ebert2018visual].
@@
Autoregressive-based world models generate each step conditioned on previous outputs, allowing them to predict sequences of arbitrary length and making them well-suited for long-horizon predictions. However, they often suffer from error accumulation over extended sequences [yang2025roboenvision] and may struggle to represent highly multi-modal distributions. In contrast, diffusion-based models generate samples through an iterative denoising process, enabling them to model complex, multi-modal distributions and produce globally coherent sequences. This iterative refinement also makes diffusion models more robust to individual prediction errors, resulting in better performance on tasks requiring long-horizon consistency or high-quality generative outputs. On the downside, diffusion models are computationally intensive and slower during inference, and adapting them to sequential prediction requires careful conditioning. Overall, autoregressive world models tend to excel in scenarios demanding speed and accurate short-term predictions. In contrast, diffusion models are more suitable for tasks such as long-horizon prediction, where maintaining global coherence is crucial.
@@
Compared with implicit and latent-space world models, video generation models provide more detailed visual predictions but at a higher computational cost, lower generation speed and sample efficiency. In addition, action predictions are only proved to be align with visual future generation [wang2025learning], as visual data contains relevant information to actions.
@@ III-B1 Flat Architecture
Most current methods adopt flat architectures [guo2025flowdreamer, ferraro2025focus, villar2025playslot, bruce2024genie, brooks2024video, xiang2024pandora, zheng2024open, zhou2024robodreamer], which face critical limitations. They lack structured representations of the environment, resulting in poor handling of multi-scale dynamics, limited long-horizon prediction, error accumulation, and reduced generalization. Specifically, in robotic manipulation, placing fragile objects requires the robot to react instantly to unexpected slips while simultaneously planning the sequence of pick-and-place actions to achieve the overall goal. Many tasks further involve long-term objectives that must be completed through sequential subgoals and temporally extended actions. For example, assembling a piece of furniture requires picking up components, aligning and attaching them correctly, and tightening screws for each part. Moreover, operating at a single level of abstraction causes small prediction errors to compound over time, degrading performance in long-horizon tasks. Finally, flat architectures fail to extract high-level abstractions, limiting transferability across tasks and environments.
III-B2 Hierarchical Architecture
Several studies have begun to explore and design hierarchical world models. In this scheme, lower-level modules handle intermediate reactions and short-term predictions, while higher-level components are responsible for long-term planning and abstraction. Lecun et al. [lecun2022path] hypothesize a hierarchical JEPA architecture, where low-level and high-level representations are learned for short- and long-term predictions, respectively. Gumbsch et al. [gumbsch2023learning] propose an RSSM-based hierarchical world model, where the low-level module captures immediate dynamics for reactive control, and the high-level module models abstract temporal patterns for strategic planning. Björck et al. [bjorck2025gr00t] introduce a dual-system architecture in which System 2 interprets the environment and task goals, while System 1 generates continuous motor commands in real time. Similarly, Wang et al. [wang2025dmwm] design a dual-level world model consisting of an RSSM-based System 1 (RSSM-S1) and a logic-integrated neural network System 2 (LINN-S2). The inter-system feedback mechanism ensures that predicted sequences comply with domain-specific logical rules: LINN-S2 constrains RSSM-S1’s predictions, while RSSM-S1 updates LINN-S2 based on new observations, enabling dynamic adaptation. Wang et al. [song2025hume] further employ System 2 for value-guided high-level planning by estimating state-action values and selecting optimal actions, while System 1 executes real-time motions via cascaded action denoising.
@@ III-C1 Dimensionality of the World
In designing world models, the dimensionality of the environment is critical in shaping a model’s ability to capture spatial structures, temporal evolution, and causal dynamics.
Some works operate purely in 2D pixel space [yang2023learning, bruce2024genie, brooks2024video, xiang2024pandora, zheng2024open, zhou2024robodreamer], capturing visual appearance and short-term dynamics but ignoring the real-world geometry. While 2D pixel-space models [yang2023learning] capturing visual appearance and short-term dynamics but lacking geometric awareness of real-world structure. This limitation motivates the development of 3D-aware architectures. To incorporate geometric understanding of the 3D world, Bu et al. [bu2024closed, ko2024learning, zhang2025combo] construct world models based on RGB-D data, while others extract richer 3D cues such as scene flow [guo2025flowdreamer], motion fields [zhi20253dflowaction] and 3D point clouds with associated language descriptions [zhen20243d], enabling more comprehensive modeling of 3D world dynamics. Additionally, Lu et al. [lu2025gwm] leverage 3D Gaussian Splatting, Diffusion Transformers, and 3D Gaussian Variational Autoencoder to extract 3D representations from RGB observations. Zhang et al. [zhang2025dreamvla] incorporate depth estimation to enhance the understanding of 3D worlds. In addition to geometric structure, temporal dynamics are incorporated to construct 4D world models that jointly capture spatial and temporal evolution. For example, Zhu et al. [team2025aether] synthesize 4D data from RGB-D videos by estimating depth and camera pose. Zhen et al. [zhen2025tesseract] leverage the pre-trained 3D VAE [kingma2013auto] to encode RGB, depth, and normal videos and sum them together. Huang et al. [huang2025enerverse] employ 4D Gaussian splatting to model spatiotemporal dynamics in robotic environments.